OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now?

OpenStudy (anonymous):

13

OpenStudy (anonymous):

Let J be Coach Johnson's age now. T will be Tim's age now. \(J=T+10\) ; (Johnson is 10 years older than Tim) \(J-3=2(T-3)\) ; (3 years ago, Johnson was twice as old as Tim). Use simultaneous equations to solve for T.